# XTS mode revisited: high hopes for key scopes?

### v1

@@ Abstract
This paper concisely summarizes the XTS block encryption mode for storage sector-based encryption applications and clarifies its limitations. In particular, we aim to provide a unified basis for much needed discussions about the newly proposed key scope change to the IEEE 1619 standard.
[…]
@@ 1 Introduction
Since the introduction of XTS [3] in 2007, many software and hardware sector-based encryption systems have widely adopted it, e.g., BitLocker, VeraCrypt, Cryptsetup, and TCG Opal self-encrypting drives. Still, it seems to be surrounded by a shroud of misunderstandings, and some of its subtler points are still sometimes contested in the developer communities. It does not help that the existing NIST XTS-AES recommendation [1] only references the paid version of the IEEE standard [3], making it probably the only NIST document that does not include a publicly available definition of the described cryptographic primitive (as already mentioned in Rogaway’s analysis [11] in 2011).
The situation is especially problematic in light of the newly proposed key scope change to the standard [4] (again behind a paywall), which would render the vast majority of implementations non-compliant. This would force vendors to request (not only) open-source projects to make significant changes to adapt to the new version. In contrast, the reasons and security implications of the proposed change are not communicated at all.
@@
  • Unify the terminology in the XTS context in a way easily understandable to different communities.
  • Describe XTS security limitations and clarify some contested points.
  • Encourage a constructive public discussion about future requirements and recommendations.
@@ 2.1 Units and sectors
The typical blocksize for modern ciphers is much smaller than the sector size (16 bytes for AES). Thus, to encrypt the whole sector, we need to use an encryption mode (such as XTS) that builds upon the original cipher. All modern real-world storage devices always use sector sizes that are multiples of 16 bytes, eliminating the need for padding during both encryption and decryption. [^2] The only legacy exception is the 520-byte sector (a 512-byte data sector with an additional 8-byte integrity field) on enterprise devices.
@@ 2.2 Notation
In this text, $E$ will denote a blockcipher (e.g., AES-128) of blocksize $n$ bits, so that for each key $K$, $E_{K}$ is a permutation on $\{0,1\}^{n}$ – the set of binary $n$-bit strings. For $s,s^{\prime}\in\{0,1\}^{n}$, we will write $s\|s^{\prime}$ and $s\oplus s^{\prime}$, for the concatenation and XOR respectively. The $s=a_{n-1}\dots a_{1}a_{0}$ can be seen as the class of the polynomial $a_{n-1}x^{n-1}+\dots+a_{1}x+a_{0}$ in the finite field $\mathbb{F}_{2^{n}}$. With this interpretation, we can multiply binary strings by elements of $\mathbb{F}_{2^{n}}$ using shift and XOR operations (though it depends on the polynomial chosen to represent $\mathbb{F}_{2^{n}}$).
@@ 3.1 XTS definition
and $\alpha:=0^{126}10$ is the class of $x$ in $\mathbb{F}_{2^{128}}:=\mathbb{F}_{2}[x]/(x^{128}+x^{7}+x^{2}+x+1)$.
Note that for a binary string $a_{127}a_{126}\dots a_{1}a_{0}$, we have
@@ 3.2 Threat model
Despite having received a lot of criticism [11, 9] after introduction, the XTS mode still has not been replaced by any other mode more than a decade later. Unfortunately, the security goals and threat models are not well-defined either.
@@ 4.1 Key scopes
The most drastic proposed change in [4] is the following requirement: “Maximum Number of 128-bit blocks in a key scope = $2^{36}$ to $2^{44}$”. For a fixed keypair $(K,K_{T})$, this translates to $S\cdot J\leq 2^{36}$ and $S\cdot J\leq 2^{44}$, respectively.
@@
There is a non-negligible probability that (1) occurs when If the number of plaintexts ($S\cdot J$) encrypted under the same keypair $(K,K_{T})$ approaches the birthday bound $2^{n/2}$.
Consequently, the scope of each keypair should be limited. The probability of (1) occurring for AES-{128,256} can be estimated by $\frac{(S\cdot J)^{2}}{2^{128}}$. Thus, as argued in Annex D.4.3 in [3], encrypting 1 TiB of data (i.e., $S\cdot J=2^{36}$) would lead to (1) occurring with probability roughly $2^{36\cdot 2-128}=2^{-53}$, whereas 256 TiB of data would correspond to roughly $2^{44\cdot 2-128}=2^{-40}$.
However, it is unclear what amount of risk is acceptable in different contexts, even though the proposed change would invalidate compliance of many current encrypted storage devices.
@@ 4.3 Distinct $K,K_{T}$
Annex C.I in [7] warns about a chosen ciphertext attack against XTS-AES, referring to Section 6 in Rogaway’s original paper [10]: “by obtaining the decryption of only one chosen ciphertext block in a given data sector, an adversary who does not know the key may be able to manipulate the ciphertext in that sector so that one or more plaintext blocks change to any desired value”. The attack crucially relies on two simultaneous conditions: $K=K_{T}$, and using a tweak corresponding to a representative of $1$ – i.e., in our case, $j=0$. There are two obvious ways to prevent the attack:
@@ 5 Discussion
The XTS encryption mode, as defined in IEEE 1619 [3], has been used for many years in many different systems. Thus, any compliance-breaking changes – such as the proposed mandatory use of key scopes in the IEEE 1619 draft [4] – could have a dramatic impact on the ecosystem. This would be amplified even more by the concept of FIPS certifications (which should impact the document [7]).
@@
  • Is there another way to strengthen XTS without the key scopes? One option would be using a blockcipher with larger $n$. As $n=256$ would already be acceptable and NIST already plans to propose a standard for “wide AES” ([8]), this option should be included in future considerations. The problem is that the XTS standard [3] has hardcoded 128-bit blocks everywhere and would need to be updated.
@@
  • How should the keypairs used for key scopes be generated? It is unclear if all the keypairs used in key scopes must be generated using an approved random number generator, or if they can be derived from a master key with an approved key derivation algorithm. The second option would simplify key scope implementation.
@@
  • How to determine which keypair to use for which sector? And how to handle device resizing? Standardizing (several variants of) answers to these might help with complexity reduction.
We propose two partial answers to the last question:
@@
  • Implement key scopes in a linear fashion: use each keypair to encrypt the next (at most) $2^{44}$ plaintexts, then switch to the next keypair. A downside is that keypairs would not be utilized uniformly: if data is stored only on a small part of the device, the keys beyond the first one are never used. Also, for device resizing, keypairs would need to be added or removed dynamically.
@@
  • Implement key scopes with rotating keypairs: use $(N\text{ mod }m)$-th keypair to encrypt the $N$-the sector, where $m$ is the total number of keypairs. This solution must define the maximal device size in advance (to not exceed the required key scope), but easily allows a dynamic resize (up to the maximal size). It also utilizes all keys more uniformly.
We believe that continuing this conversation publicly would be especially relevant not only for the hardware world, where vendors would like to reuse existing accelerated implementations. From a long-term perspective, it might be more beneficial to switch to a different (wide) encryption mode (as suggested in [11]) if length-preserving ciphertext is required. Or if the storage device provides space for authentication tags, authenticated encryption would be a strong candidate as well.

### v2

@@ Abstract
This paper concisely summarizes the XTS block encryption mode for storage sector-based encryption applications and clarifies its limitations. In particular, we aim to provide a unified basis for constructive discussions about the newly introduced key scope change to the IEEE 1619 standard. We also reflect on wide modes that could replace XTS in the future.
This work is licensed under a “CC BY ” license.
@@ 1 Introduction
Since the introduction of XTS [10] in 2007, many software and hardware sector-based encryption systems have widely adopted it, e.g., BitLocker [14], VeraCrypt [2], Cryptsetup [1], and TCG Opal [8] self-encrypting drives. Still, it seems to be surrounded by a shroud of misunderstandings, and some of its subtler points are still sometimes contested in the developer communities. It does not help that the existing NIST XTS-AES recommendation [7] only references the paid version of the IEEE standard [10], making it probably the only NIST document that does not include a publicly available definition of the described cryptographic primitive (as already mentioned in Rogaway’s analysis [20] in 2011).
The situation is especially problematic in light of the newly introduced key scope change to the IEEE 1619-2025 standard [12] (again behind a paywall), which would render the vast majority of implementations non-compliant. This would force vendors to request (not only) open-source projects to make significant changes to adapt to the new version. In contrast, the reasons and security implications of the change are not communicated at all.
@@
  • Unify the terminology in the XTS context in a way easily understandable to both theoreticians and practicioners (Section 2 and 3).
  • Describe XTS security limitations (namely key scopes, maximal sector size, and distinct keys) and clarify some contested points (Section 4).
  • Encourage a constructive public discussion to influence future requirements and recommendations (Section 5).
@@ 2.1 Units and sectors
The typical blocksize for modern blockciphers is much smaller than the sector size (16 bytes for AES). Thus, to encrypt the whole sector, we need to use an encryption mode (such as XTS) that builds upon the original cipher. All modern real-world storage devices always use sector sizes that are multiples of 16 bytes, eliminating the need for padding during both encryption and decryption. [^2] The only legacy exception is the 520-byte sector (a 512-byte data sector with an additional 8-byte integrity field) on enterprise devices.
@@ 2.2 Notation
In this text, $E$ will denote a blockcipher (e.g., AES-128) of blocksize $n$ bits, so that for each key $K$, $E_{K}$ is a permutation on $\{0,1\}^{n}$ – the set of binary $n$-bit strings. For $s,s^{\prime}\in\{0,1\}^{n}$, we will write $s\|s^{\prime}$ and $s\oplus s^{\prime}$, for the concatenation and XOR operations, respectively. A binary string with bits $a_{n-1},\dots,a_{1},a_{0}$ corresponds to the element $a_{n-1}x^{n-1}+\dots+a_{1}x+a_{0}]$ of the finite field $\mathbb{F}_{2^{n}}$. Thus, we can multiply binary strings by elements of $\mathbb{F}_{2^{n}}$ using shift and XOR operations (though it depends on the polynomial chosen to represent $\mathbb{F}_{2^{n}}$).
@@ 3.1 XTS definition
and $\alpha$ is the element $[x]$ in $\mathbb{F}_{2^{128}}:=\mathbb{F}_{2}[x]/(x^{128}+x^{7}+x^{2}+x+1)$.
Conceptually, the multiplication by $\alpha$ corresponds to a left shift with overflow. More explicitly, for a binary string $s=a_{127}a_{126}\dots a_{1}a_{0}$ (this is how we denote concatenation of bits), we have
@@ 3.2 Threat model
Despite having received a lot of criticism [20, 18], the XTS mode has not been replaced by any other mode more than a decade later. Unfortunately, its security goals and threat models are still not well-defined.
@@ 4.1 Key scopes
The most drastic change in the IEEE 1619-2025 standard [12] is the following requirement: “Maximum Number of 128-bit blocks in a key scope = $2^{36}$ to $2^{44}$”. For a fixed XTS key [^5] By an XTS key, we really mean a concatenated pair of symmetric keys; this is standard terminology. $(K,K_{T})$, this translates to $S\cdot J\leq 2^{36}$ and $S\cdot J\leq 2^{44}$, respectively.
@@
There is a non-negligible probability that (1) occurs when the number of plaintexts ($S\cdot J$) encrypted under the same XTS key $(K,K_{T})$ approaches the birthday bound $2^{n/2}$.
Consequently, the scope of each key should be limited. The probability of (1) occurring for AES-{128,256} can be estimated by $\frac{(S\cdot J)^{2}}{2^{128}}$. Thus, encrypting 1 TiB of data (i.e., $S\cdot J=2^{36}$) would lead to (1) occurring with probability roughly $2^{36\cdot 2-128}=2^{-56}$, whereas 256 TiB of data would correspond to roughly $2^{44\cdot 2-128}=2^{-40}$. This agrees with Annex D.4.3 in [10] up to a small constant. [^6] Annex D.4.3 in [10] mistakenly uses a general security guarantee $9.5\frac{q^{2}}{2^{n}}$ from [19] (valid for XEX, not XTS), but in fact, the factor $9.5$ can be dropped if AES is secure [13, 3]. However, it is unclear what amount of risk is acceptable in different contexts, [^7] We discussed this issue both in an IEEE SISWG meeting on April 25, 2025 and associated private communication, and got a confirmation that the analysis 4.1 is the one motivating the key scopes. The goal was to satisfy security targets set by NIST (i.e., $2^{-53}$ and $2^{-37}$, respectively, should be good enough), but NIST never explained the choice of the specific constants, just like with the maximal sector size in Section 4.2. even though the change would invalidate compliance of many current encrypted storage devices.
@@ 4.3 Distinct $K,K_{T}$
Annex C.I in [16] warns about a chosen ciphertext attack against XTS-AES, referring to Section 6 in Rogaway’s original paper [19]: “by obtaining the decryption of only one chosen ciphertext block in a given data sector, an adversary who does not know the key may be able to manipulate the ciphertext in that sector so that one or more plaintext blocks change to any desired value”. The attack crucially relies on two simultaneous conditions: $K=K_{T}$, and starting the indexation at $j=0$ (where the tweak is unchanged, as it is multiplied by $\alpha^{0}$). There are two obvious ways to prevent the attack:
@@ 5 Discussion
The XTS encryption mode, as defined in IEEE 1619 [10], has been used for many years in many different systems. Thus, any compliance-breaking changes – such as the mandatory use of key scopes in the IEEE 1619 standard [12] – could have a dramatic impact on the ecosystem. This would be amplified even more by the concept of FIPS certifications (which should impact the document [16]).
@@
  • Is there another way to strengthen XTS without the key scopes? One option would be using a blockcipher with larger $n$. As $n=256$ would already be acceptable, and NIST already plans to propose a standard for “wide AES” ([17]), this option should be included in future considerations. The problem is that the XTS standard [10] has hardcoded 128-bit blocks everywhere and would need to be updated.
@@
  • How should the XTS keys used for key scopes be generated? It is unclear if all the keys used in key scopes must be generated using an approved random number generator, or if they can be derived from a master key with an approved key derivation algorithm. The second option would simplify key scope implementation.
@@
  • How to determine which XTS key to use for which sector? And how to handle device resizing? Standardizing (several variants of) answers to these might help with complexity reduction.
Currently, XTS-AES uses a single XTS key for the whole device (for all addressable sectors). Introducing key scopes can lead to incompatible variants of XTS-AES implementations until we specify which XTS key corresponds to a given sector. We propose two partial answers to this problem:
@@
  • Implement key scopes in a linear fashion: use each XTS key to encrypt the next (at most) $2^{44}$ plaintexts, then switch to the next key. A downside is that keys would not be utilized uniformly: if data is stored only on a small part of the device, the keys beyond the first one are never used. Also, for device resizing, keys would need to be added or removed dynamically.
@@
  • Implement key scopes with rotating XTS keys: use $(N\text{ mod }m)$-th XTS key to encrypt the $N$-the sector, where $m$ is the total number of XTS keys. This solution must define the maximal device size in advance (to not exceed the required key scope), but easily allows a dynamic resize (up to the maximal size). It also utilizes all keys more uniformly.
Beyond XTS mode
We would like to stress that XTS comes with serious downsides, just like other classical modes such as ECB, CBC, CFB, OFB, and CTR. Namely, they fail to provide full diffusion (i.e., each bit of the ciphertext depending on each bit of the plaintext) if the data to be encrypted exceeds a few blocks [6].
For a long-term use of a length-preserving mode (i.e., disqualifying authenticated encryption), it seems much more fruitful to switch to a wide encryption mode suited for disk encryption (as suggested in [20]). The feasible candidates might include:
  • EME2 is based on EME, designed in 2003 and later standardized by IEEE 1619.2 [11]. The mode was never widely adopted, partially because of a questionable status of the related patent (now abandoned [^8] https://lore.kernel.org/dm-crypt/kf5lug\$8p0\$1@ger.gmane.org/).
  • Adiantum [4] was developed by Google for low-end systems that do not provide AES hardware acceleration. It is designed to be fast and has low power requirements (suitable for mobile devices).
  • HCTR2 [5] (also by Google) builds upon the CTR mode and seems like a performant and conservative option.
  • bbb-ddd-AES [6] is a new mode built over the docked-double-decker construction [9]. If there is an upper bound on the number of times each tweak can be reused (sometimes satisfied by SSD hardware due to limited lifetime and wear-leveling), it can provide beyond-birthday bound security. The double-decker construction is a flexible framework that allows building wide encryption modes with better characteristics than the older modes mentioned above. As the proposed modes are very recent, there is no independent security analysis, and – unlike for the above candidates – no public reference implementation.
The pragmatic solution is perhaps to keep XTS limits in a sustainable state, as many legacy systems will be using it in the next years. For long-term improvement of disk encryption, we believe it would be worthwile to focus on a newer construction, like the mentioned double-decker framework.